[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Apr 29, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Fri Apr 29 16:46:52 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The problem with the shiner N-AWIPS workstation that did not reboot  
properly was investigated further, and that system does not have the  
procedure to check for missing video drivers. We will work with NCO to get  
the implemented on all the operational N-AWIPS systems.
HSU:
None.
Other items:
There was a problem with the AC systems in the perimeter offices early this  
morning, which was resolved by about 8:30 AM. Patty let the facilities  
manager know so the cause of the problem can be tracked down.
NHC RFCs:
None.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160429/413a6fbe/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list